OpenStudy (anonymous):

square root of 45 and 125?

OpenStudy (immanuelv):

\[\sqrt{45}=3\sqrt{5}\]

OpenStudy (immanuelv):

\[\sqrt{125}=5\sqrt{5}\]

OpenStudy (anonymous):

thanks!